OpenStudy (anonymous):

solve the boundary value problem if possible. \[y''-3y'+2y=0\] y(0)=1 y(3)=0 auxiliary equation \[r^2-3r+2=0\] \[(r-2)(r-1)=0\] \[r_1=2\] \[r_2=1\] \[y(x)=c_1e^{2x}+c_2e^{x}\] \[y(0)=c_1e^{2(0)}+c_2e^0=1\] \[y(0)=c_1+c_2=1\] \[y(3)=c_1e^6+c_2e^3=0\] \[e^6-e^6c_2+c_2e^3=0\] \[-c_2+c_2\frac1{e^3}=0\] \[\frac1{e^3}=1\]

OpenStudy (anonymous):

\[1-c_2+c_2\frac1{e^3}=0\]

OpenStudy (anonymous):

everything looked fine until this y(0)=c1+c2=1 y(3)=c1e^6+c2e^3=0 c1+c2=1 c1*e^6+c2e^3=0 2 equations 2 unknowns solve for c1 and c2 e^6 and e^3 are both constants

OpenStudy (anonymous):

one moment...

OpenStudy (anonymous):

e^6=403.428555 e^3=20.08553692 c1+c2=1 403.429*c1+20.086*c2=0 c1=1-c2 403.429(1-c2)+20.086c2 =0 403.429-403.429c2+20.086c2=0 403.329=(403.429-20.086)c2

OpenStudy (anonymous):

\[c_2=\frac1{1-e^{\frac13}}\]

OpenStudy (anonymous):

\[c_1=1-\frac1{1-e^{\frac13}}\]

OpenStudy (anonymous):

\[c2=\frac{e^6}{e^6-e^3}\] if you divided top and bottom by e^6 wont you get \[c2=\frac{1}{1-e^{-3}}\] or did i get my math wrong @MathSofiya

OpenStudy (anonymous):

oh just realized you were the one asking the question

OpenStudy (anonymous):

If the statement below is correct \[e^6-e^6c_2+c_2e^3=0\] then wouldn't we get \[1-c_2+c_2\frac1{e^3}=0\] since \[\frac{e^3}{e^6}=\frac1{e^3}\]

OpenStudy (anonymous):

where does the e^1/2 come from?

OpenStudy (anonymous):

meh ill re do it using actual exponents

OpenStudy (anonymous):

what do you mean e^.5?

OpenStudy (anonymous):

i mean you answer for c2

OpenStudy (anonymous):

\[1-c_2+c_2\frac1{e^3}=0\] \[1=c_2-c_2\frac1{e^3}\] \[1=c_2(1-\frac1{e^3})\]

OpenStudy (anonymous):

then solve for c_2 \[c_2=\frac1{1-e^{\frac13}}\]

OpenStudy (anonymous):

\[c_1+c_2=1\] \[c_1e^6+c_2e^3=0\] \[c_1=1-c_2\] \[(1-c_2)e^6+c_2e^3=0\] \[e^6-c_2e^6+c_2e^3=0\] \[e^6=c_2(e^6-e^3)\] \[c_2=\frac{e^6}{e^6-e^3}\] \[c_2=\frac{\frac{e^6}{e^6}}{\frac{e^6-e^3}{e^6}}\] \[c_2=\frac{1}{1-\frac{1}{e^3}}\] \[\frac{1}{e^3}=e^{-3}\] \[c_2=\frac{1}{1-e^{-3}}\]

OpenStudy (anonymous):

so I was right?

OpenStudy (anonymous):

c_1=1-\frac1{1-e^{\frac13}}

OpenStudy (anonymous):

\[c_1=1-\frac1{1-e^{\frac13}}\]

OpenStudy (anonymous):

and \[c_2=\frac1{1-e^{\frac13}}\]

OpenStudy (anonymous):

using your last step before you solved for c2 \[1=c_2(1−\frac{1}{e^3})\] \[1=c_2(1−e^{-3})\] \[c_2\frac{1}{(1−e^{-3})}\]

OpenStudy (anonymous):

e^-3 is very different from e^(1/2)

OpenStudy (anonymous):

\[e^{\frac{1}{2}}=\sqrt{e}\]

OpenStudy (anonymous):

that's when I made a mistake and I corrected it My mistake was \[-c_2+c_2\frac1{e^3}=0\] and it should have been \[1-c_2+c_2\frac1{e^3}=0\]

OpenStudy (anonymous):

your final solution shouldnt have e^(1/2) assuming i didnt read the wrong problem

OpenStudy (anonymous):

scroll up I a little I have my solutions for c_1 and c_2 about 5 posts up

OpenStudy (anonymous):

i saw it and im saying its incorrect.... the 1/2 should be -3

OpenStudy (anonymous):

\[c_1=1-\frac1{1-e^{\frac13}}\] \[c_2=\frac1{1-e^{\frac13}}\] I've posted them also 24 minutes ago

OpenStudy (anonymous):

the only way you get 1/2 from 1/e^3 is if you had them to the -1/6th power

OpenStudy (anonymous):

where are you getting 1/2 from?

OpenStudy (anonymous):

sorry i mean it shouldnt be e^1/3

OpenStudy (anonymous):

the 3 looks like a 2 too small

OpenStudy (anonymous):

oh I see...why not though?

OpenStudy (anonymous):

@satellite73 what do you think?

OpenStudy (anonymous):

\[\frac{1}{e^3}=e^{-3}\] \[e^{\frac{1}{3}}=\sqrt[3]{e}\]

OpenStudy (anonymous):

\[1−c_2+c_2\frac{1}{e^3}=0\] \[1=c_2-c_2\frac{1}{e^3}\] \[1=c_2(1-\frac{1}{e^3})\] \[c_2=\frac{1}{1-\frac{1}{e^3}}\] \[c_2=\frac{1}{1- e^{-3}}\]

OpenStudy (anonymous):

are we agreeing? I posted the question again to the left since this post is getting a little long

OpenStudy (anonymous):

\[1−c_2+c_2\frac{1}{e^3}=0\] \[1−c_2+c_2e^{-3}=0\]

OpenStudy (anonymous):

I think we come to the same conclusion, I think the main confusion was the 1/3 looking like a 1/2

OpenStudy (anonymous):

ill just say it c2 is equal to 1 over (1 minus e to the negative 3rd power)

OpenStudy (anonymous):

i keep seeing your answer as c2 is equal to 1 over (1 minus e to the one-third power)

OpenStudy (anonymous):

what's the difference?

OpenStudy (anonymous):

e to the one third is not equal to e to the negative third whenever you have fractions in exponents, it means the root of x^(1/2)= the square root of x x^(1/3)= the third root of x x^-1 = 1 over x x^-2= 1 over x^2 if x=2 x^(1/2)=2 x^-2= 1/4 2 is not equal to 1/4